slope is -6/7 and it passes through (14,-3)
Answer by nerdybill(7384)

slope is -6/7 and it passes through (14,-3)
.
Plug given info into "point-slope" form:
y - y1 = m(x - x1)
y - (-3) = (-6/7)(x - 14)
y + 3 = (-6/7)(x - 14)
y + 3 = (-6/7)x + 12
y = (-6/7)x + 9